2[4 – 2(3 – x)] – 1 = 4[2(4x – 3) + 7] – 25 can you help me solve this and show me how to work it out ? I understand that there are different ways. Can someone please show me some ? Thanks

Answer by checkley71(8403)

THERE IS ONLY 1 WAY TO SOLVE THESE TYPE OF PROBLEMS.
WORK FROM WITHIN THE ( ) FIRST, THEN MULTIPLY/DIVIDE THEN ADD/SUBTRACT.
2[4-2(3-X)-1=4[2(4X-3)+7]-25
2[4-6+2X]-1=4[8X-6+7]-25
8-12+4X-1=32X-24+28-25
4X-32X=-24+28-25-8+12
-28X=-17
X=-17/-28
X=17/28 ANSWER.